OpenStudy (anonymous):

The length of a diagonal of a square is 7 inches. Find the perimeter of the square. A. 25 inches B. 12 inches C. 19.6 inches D. 28 inches

OpenStudy (anonymous):

7 is the hypotenuse in the following isosceles right triangle: So, a^2 + b^2 = c^2 -> 2a^2 = 7^2 -> a = 7 / [sqrt(2)] P = 4a = 4 times 7 / [sqrt(2)] = (14)[sqrt(2)]

OpenStudy (anonymous):

So, 14 x sqrt of 2?

OpenStudy (anonymous):

That triangle is part of: Yes, 14 times sqrt(2)

OpenStudy (anonymous):

19

OpenStudy (anonymous):

Yes! (rounded) It's actually 19.798989873223330683223642138936

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

You're getting good at these!